OpenStudy (anonymous):

Use the “difference of squares” rule to factor the following expression: 64z⁴-81a⁴

OpenStudy (anonymous):

\[64z^4-81a^4=(8z^2)^2-(9a^2)^2=(8z^2-9a^2)(8z^2+9a^2)\] \[(8z^2-9a^2)(8z^2+9a^2)=((2\sqrt{2}z)^2-(3a)^2)(8z^2+9a^2)=(2\sqrt{2}z-3a)(2\sqrt{2}z+3a)(8z^2+9a^2)\]